OpenStudy (anonymous):

The area of the region bounded by the lines x=0,x=2, and y=0 and the curve y=e^(x/2) is?

OpenStudy (anonymous):

\[required ~area=\int\limits_{0}^{2} y~ dx=\int\limits_{0}^{2} e ^{\frac{ x }{ 2 }}dx\] putx/2=t x=2t dx=2dt when~x=0,t=0 when x=2,t=2/2=1 \[required area=2\int\limits_{0}^{1}e ^{t}~dt\] i think now you can complete it.